OpenStudy (anonymous):

Factor the trinomial: x2 + 12x + 32 (x - 2)(x + 16) (x + 16)(x + 16) (x + 4)(x + 8) prime could someone show me how to do this?

OpenStudy (anonymous):

x^2

OpenStudy (anonymous):

(x+4) (x+8)

OpenStudy (anonymous):

x^2 +4x+8x+4*8= x^2+12x+32

OpenStudy (anonymous):

do u see how this works?

OpenStudy (anonymous):

kinda of but thank you

OpenStudy (anonymous):

since x is squared, each parenthesis will start with x. and since both symbols are +, so will both of the parenthesis pairs. Nverleur is right with the answer because 4+8=12 (the middle number) and 4 times 8=32. so there you go. hope this helped.

OpenStudy (anonymous):